# Spectral selectors and contact orderability

### v1

@@ 1.1. Historical background
The abundance of interactions between Hofer geometry on the group of Hamiltonian diffeomorphisms and symplectic geometry [43, 46, 51, 61] led Eliashberg and Polterovich in 2000 [30] to introduce the notion that will be later called orderability on the universal cover $\widetilde{\mathcal{G}}$ of the group of contactomorphisms isotopic to the identity. To do so they define a bi-invariant binary relation $\preceq$ on this group: $\varphi\preceq\psi$ if there exists a non-negative contact Hamiltonian generating an isotopy from $\varphi$ to $\psi$. They show that for some contact manifolds this binary relation turns out to be a partial order: we now say that the group $\widetilde{\mathcal{G}}$ is orderable in this case. Some years later together with Kim [29] they show how orderability can be used to detect some squeezing and non-squeezing phenomena. They show in particular that there really is a dichotomy: for some contact manifolds such as the standard contact sphere $\widetilde{\mathcal{G}}$ is unorderable.
The study of orderability naturally extends to the group of contactomorphisms isotopic to the identity $\mathcal{G}$ and to the isotopy class $\mathcal{L}$ of a closed Legendrian submanifold $\Lambda_{*}$ as well as its universal cover $\widetilde{\mathcal{L}}$. The notion of orderability has been quite influential in the study of contact topology: we refer to our paragraph Examples 2.10 in Section 2.4 for an account on the study of the dichotomy orderable/unorderable space throughout the last two decades. Nevertheless, rather little is known about the properties that are specifically implied by orderability. One of the most significant achievements in that direction is the discovery of Albers-Fuchs-Merry that the unorderability of $\mathcal{G}(M,\xi)$ implies the Weinstein conjecture in $(M,\xi)$ [2]. The goal of this article is to establish an equivalence between orderability and the existence of (hopefully spectral) contact selectors.
@@
    (triangular inequality) $c_{+}(\varphi\psi)\leq c_{+}(\varphi)+c_{+}(\psi)$,
@@ 1.2. Conventions
Whenever no precision regarding regularity is given, maps and manifolds considered are $C^{\infty}$-smooth. Every contact manifold $(M,\xi)$ considered are assumed to be connected and cooriented (i.e. $TM/\xi$ is an oriented line bundle). We write that a contact form $\alpha$ is supporting $\xi$ if $\ker\alpha=\xi$ and $\alpha$ respects the coorientation. A contactomorphism of $(M,\xi)$ will always preserve the coorientation. Given a contact manifold $(M,\xi)$ and a closed Legendrian submanifold $\Lambda_{*}\subset M$, one denotes $\mathcal{L}(\Lambda_{*})$ the isotopy class of $\Lambda_{*}$ and ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ its universal cover. One denotes $\mathcal{G}(M)$ the group $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ of compactly supported contactomorphisms of $(M,\xi)$ isotopic to $\mathrm{id}$ through compactly supported contactomorphisms while $\widetilde{\mathcal{G}}(M)$ will denote its universal cover.
@@
On $O$ being either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, we write $x\preceq y$, or equivalently $y\succeq x$, if there exists a non-negative isotopy from $x$ to $y$. $O$ is called orderable if and only if $\preceq$ defines a partial order (see Section 2.4 for details). In particular, the orderability of $\mathcal{L}$ (resp. $\mathcal{G}$) implies the orderability of $\widetilde{\mathcal{L}}$ (resp. $\widetilde{\mathcal{G}}$). We refer to Examples 2.10 for a list of known orderable spaces.
@@ 1.3. Order spectral selectors
Let $\Lambda_{*}$ be a closed Legendrian submanifold of $(M,\xi)$ such that $\mathcal{L}(\Lambda_{*})$ (resp. ~$\widetilde{\mathcal{L}}(\Lambda_{*})$) is orderable and let $\alpha$ be a complete contact form supporting $\xi$. The selectors ±$\ell_{\pm}^{\alpha}$ are real-valued and satisfy the following properties for every $\Lambda_{0},\Lambda_{1},\Lambda_{2}\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$),
@@
    (compatibility) ±$\ell_{\pm}^{\alpha}(\varphi(\Lambda_{1}),\varphi(\Lambda_{0}))=\ell_{\pm}^{\varphi^{*}\alpha}(\Lambda_{1},\Lambda_{0})$, for every $\varphi\in\mathrm{Cont}(M,\xi)$ (resp. $\widetilde{\mathrm{Cont}}(M,\xi)$),
@@
    (non-degeneracy) $\ell_{+}^{\alpha}(\Lambda_{1},\Lambda_{0})=\ell_{-}^{\alpha}(\Lambda_{1},\Lambda_{0})=t$ for some $t\in\mathbb{R}$ implies $\Lambda_{1}=\phi_{t}^{\alpha}\Lambda_{0}$ (resp. it only implies the equality $\Pi\Lambda_{1}=\phi_{t}^{\alpha}\Pi\Lambda_{0}$ in $\mathcal{L}$).
@@
    (compatibility) $c_{\pm}^{\alpha}(\varphi\psi\varphi^{-1})=c_{\pm}^{\varphi^{*}\alpha}(\psi)$, which extends to every $\varphi\in\mathrm{Cont}(M,\xi)$ (resp. $\widetilde{\mathrm{Cont}}(M,\xi)$),
@@
    (non-degeneracy) $c_{-}^{\alpha}(\psi)=c_{+}^{\alpha}(\psi)=t$ for some $t\in\mathbb{R}$ implies $\psi=\phi^{\alpha}_{t}$ (resp. it only implies the equality $\Pi\psi=\phi^{\alpha}_{t}$ in $\mathcal{G}$).
@@
We emphasize the fact that we do not know that the $c_{\pm}^{\alpha}(\psi)$’s do actually select spectral values of $\psi$. Nevertheless, we strongly believe it should be the case. The following direct corollary states that orderability is equivalent to the existence of selectors in a rather weak sense (a similar statement for spaces $\mathcal{L}$ and $\widetilde{\mathcal{L}}$ also holds).
In contrast with their symplectic counterparts, these selectors are not actually invariant: identities $\ell(\varphi\Lambda_{1},\varphi\Lambda_{0})=\ell(\Lambda_{1},\Lambda_{0})$ and $c(\varphi\psi\varphi^{-1})=c(\psi)$ are not verified for every contactomorphism $\varphi$ (resp. lift in the universal cover), $\Lambda_{0},\Lambda_{1}$ and $\psi$ being fixed. In that sense, the $\alpha$-selectors are only invariant under $\alpha$-strict contactomorphisms: ±$\ell^{\alpha}_{\pm}(\varphi\Lambda_{1},\varphi\Lambda_{0})=\ell^{\alpha}_{\pm}(\Lambda_{1},\Lambda_{0})$ and $c^{\alpha}_{\pm}(\varphi\psi\varphi^{-1})=c^{\alpha}_{\pm}(\psi)$ when $\varphi^{*}\alpha=\alpha$. Of course, this is compatible with the fact that the $\alpha$-spectrum is only invariant under strict contactomorphisms: as in general $\varphi^{-1}\phi^{\alpha}_{t}\varphi=\phi^{\varphi^{*}\alpha}_{t}$ for all $t\in\mathbb{R}$, one has
@@
Nonetheless, the sign of these selectors is invariant (see Lemmata 3.2 and 3.8)
@@ 1.4. Applications to the existence of Reeb chords and translated points
In particular, if $(M,\xi)$ is a closed contact manifold such that ~$\widetilde{\mathcal{L}}(\Delta\times\{0\})$ as defined in Example 2.10.6 is orderable, any contactomorphism of $M$ isotopic to $\mathrm{id}$ has an $\alpha$-translated point (and at least two distinct couples $(p,t)$ where $p$ is a translated point of time-shift $t$), given any $\alpha$ supporting $\xi$.
@@
In its original setting, the Arnold chord conjecture asked about the existence of a non-trivial Reeb chord the endpoints of which lie on a same Legendrian circle of the standard contact 3-sphere, given any Legendrian circle [9, §8]. In this setting Mohnke proved a generalization to closed contact manifolds arising as the boundary of subcritical Stein manifolds (Legendrian circles are replaced by closed Legendrian submanifolds) [57]. Hutching and Taubes extended it to any closed contact 3-manifolds [45]. Very recent works still discuss generalized forms of the original conjecture: Chantraine proved a version for a specific kind of fillable contact manifolds with Lagrangian slices taking the place of Legendrian submanifolds [17], the case of Legendrian submanifolds lying in $P\times\mathbb{R}$, for a Liouville manifold $P$, has also received a lot of attention (see [49] and references therein).
@@
Since the discovery by Hofer of a bi-invariant metric on the group of Hamiltonian diffeomorphisms [43], multiple attempts to obtain a contact counterpart where made. On the one hand, the obvious generalization of the Hofer metric, often referred to as the Shelukhin-Hofer metric, is not bi-invariant. On the other hand, one can define interesting bi-invariant metrics on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ but they are discrete, as was discovered by Colin-Sandon [26] and Fraser-Polterovich-Rosen [35]. Our spectral selectors allow us to define a contact analogue of the spectral metric discovered by Viterbo in [73]. This metric has applications to the study of both Hofer type contact metrics and discrete bi-invariant metrics of Colin-Sandon and Fraser-Polterovich-Rosen on the questions of non-degeneracy, unboundedness, metric-equivalence and geodesics. We directly refer to Section 4 for details.
@@ 2.1. Legendrian and contact isotopies and their Hamiltonian maps
Throughout this section $I\subset\mathbb{R}$ will always be an interval containing 00. Let $(M,\xi)$ be a cooriented contact manifold. A contact form $\alpha$ supporting $\xi$ (i.e. $\ker\alpha=\xi$) is called complete if its Reeb vector field $R^{\alpha}$ induces a complete Reeb flow. We recall that the Reeb vector field $R^{\alpha}$ is uniquely defined by the equations $\alpha(R^{\alpha})\equiv 1$ and $\iota_{R^{\alpha}}\mathrm{d}\alpha=0$. Let us fix a complete contact form $\alpha$ supporting $\xi$.
Let us recall that for any contact isotopy (that is a smooth isotopy of contactomorphisms) $(\varphi_{t})_{t\in I}$ with $\varphi_{0}=\mathrm{id}$, there exists a unique smooth map $H:I\times M\to\mathbb{R}$ related to the time-dependent contact vector field $(X_{t})_{t\in I}$ generating the isotopy by
@@
where $R^{\alpha}$ denotes the Reeb vector field of $\alpha$ and $H_{t}:M\to\mathbb{R}$ is induced by the restriction of $H$ to $\{t\}\times M$. (see e.g. [36, §2.3]). The smooth map $H$ is called the $\alpha$-contact Hamiltonian map of $(\varphi_{t})$ (Hamiltonian map for short). Conversely, to any smooth map $H:I\times M\to\mathbb{R}$ is associated a unique time-dependent contact vector field $(X_{t})$ by the system (2). When the contact vector field is complete (e.g. when the differential of $H$ is compactly supported), we say that $H$ generates the contact isotopy obtained by integrating $(X_{t})$.
@@
As $j_{t}^{*}\alpha=0$ for all $t\in I$, one can deduce that $(H_{t})$ only depends on the isotopy $(\Lambda_{t})$ and not the specific choice of parametrization $j$. The family $(H_{t})$ is called the $\alpha$-contact Hamiltonian map of $(\Lambda_{t})$. The contact Hamiltonian map can be seen as a smooth map $H:N\to\mathbb{R}$ defined on the submanifold $N:=\bigcup_{t}\{t\}\times\Lambda_{t}$ of $I\times M$. If $\Lambda_{t}=\varphi_{t}(\Lambda_{0})$ for some contact isotopy $(\varphi_{t})$ generated by the Hamiltonian $K$, the Hamiltonian of $(\Lambda_{t})$ is the restriction of $K$ to $N$.
@@
Let $I\subset\mathbb{R}$ be an interval containing 00 and let $(\Lambda_{t})_{t\in I}$ be a Legendrian isotopy on $(M,\ker\alpha)$, the $\alpha$-contact Hamiltonian map of which is $H:\bigcup_{t}\{t\}\times\Lambda_{t}\to\mathbb{R}$. Let $K:I\times M\to\mathbb{R}$ be an $\alpha$-contact Hamiltonian smoothly extending $H$ and generating a contact isotopy $(\varphi_{t})$, then $\varphi_{t}(\Lambda_{0})=\Lambda_{t}$ for all $t\in I$.
Before proving Lemma 2.1, let us recall the link between Legendrian isotopies and the Hamilton-Jacobi equation. Let $(\Lambda_{t})$ be a Legendrian isotopy. According to the Legendrian Weinstein neighborhood theorem, there exists an open subset $W\subset M$ containing $\Lambda_{0}$ that is contactomorphic to a neighborhood of the zero-section of the 1-jet space $J^{1}\Lambda_{0}:=T^{*}\Lambda_{0}\times\mathbb{R}$, through this identification $\Lambda_{0}$ is sent to the zero-section. Moreover, our chosen contact form $\alpha$ is sent to the standard contact form $\alpha_{0}:=\mathrm{d}z-\lambda$, $z$ being the $\mathbb{R}$-coordinate and $\lambda$ being the pull-back of the tautological form of $T^{*}\Lambda_{0}$. With this local identification, for $t\in I$ close to 00, $\Lambda_{t}$ is a Legendrian graph above the zero section $\Lambda_{0}$, so it is the $1$-jet of a smooth map $f_{t}:\Lambda_{0}\to\mathbb{R}$,
@@
for all $q\in\Lambda_{0}$ and $t\in I$ close to 00 (we have used that $\alpha$ is identified to $\alpha_{0}$ in $W$).
@@
By connectivity of the interval $I$, it is enough to prove this statement for $t\in I$ close enough to 00. By considering a Weinstein neighborhood of $\Lambda_{0}$ as above, one can then identify $\Lambda_{t}$ and $\varphi_{t}(\Lambda_{0})$ to the $1$-jet of the respective maps $f_{t}:\Lambda_{0}\to\mathbb{R}$ and $g_{t}:\Lambda_{0}\to\mathbb{R}$, for all $t\in I$, shrinking $I$ if necessary. Let us remark that the contact Hamiltonian map of $(\varphi_{t}(\Lambda_{0}))$ is the restriction of $K$ to $\bigcup_{t}\{t\}\times\varphi_{t}(\Lambda_{0})$. Therefore, by the above discussion, both families of maps $(f_{t})$ and $(g_{t})$ satisfy the same Hamilton-Jacobi equation
@@
with the initial condition $u_{0}\equiv 0$. By unicity of the smooth solutions of the Hamilton-Jacobi equation in a small interval of time (see e.g. [52, Exercice 3.5.17]), $f_{t}=g_{t}$ for all $t\in I$ and $\Lambda_{t}=\varphi_{t}(\Lambda_{0})$. ∎
@@
Let $(M,\ker\alpha)$ be a contact manifold endowed with a complete contact form $\alpha$. Let $I\subset\mathbb{R}$ be an interval containing 00 and let $(\Lambda_{t})_{t\in I}$ be a (closed) Legendrian isotopy on $(M,\ker\alpha)$, the $\alpha$-contact Hamiltonian map of which is $H$. Then there exist contact Hamiltonians $K,G:I\times M\to\mathbb{R}$ such that $\inf K=\inf H$ and $\sup G=\sup H$, the respective contact flows of which $(\varphi_{t})$ and $(\psi_{t})$ satisfy $\varphi_{t}(\Lambda_{0})=\psi_{t}(\Lambda_{0})=\Lambda_{t}$ for all $t\in I$.
@@
2.2. The $C^{1}$-topology of Legendrian and contact isotopy classes

### v2

@@ 1.1. Historical background
The abundance of interactions between Hofer geometry on the group of Hamiltonian diffeomorphisms and symplectic geometry [44, 47, 52, 62] led Eliashberg and Polterovich in 2000 [31] to introduce the notion that will be later called orderability on the universal cover $\widetilde{\mathcal{G}}$ of the group of contactomorphisms isotopic to the identity. To do so they define a bi-invariant binary relation $\preceq$ on this group: $\varphi\preceq\psi$ if there exists a non-negative contact Hamiltonian generating an isotopy from $\varphi$ to $\psi$. They show that for some contact manifolds this binary relation turns out to be a partial order: we now say that the group $\widetilde{\mathcal{G}}$ is orderable in this case (see Section 2.4 for more details). Some years later together with Kim [30] they show how orderability can be used to detect some squeezing and non-squeezing phenomena. They show in particular that there really is a dichotomy: for some contact manifolds such as the standard contact sphere $\widetilde{\mathcal{G}}$ is unorderable.
The study of orderability naturally extends to the group of contactomorphisms isotopic to the identity $\mathcal{G}$ and to the isotopy class $\mathcal{L}$ of a closed Legendrian submanifold $\Lambda_{*}$, i.e. the set of Legendrians that are Legendrian isotopic to $\Lambda_{*}$, as well as its universal cover $\widetilde{\mathcal{L}}$. The notion of orderability has been quite influential in the study of contact topology: we refer to our paragraph Examples 2.14 in Section 2.4 for an account on the study of the dichotomy orderable/unorderable space throughout the last two decades. Nevertheless, rather little is known about the properties that are specifically implied by orderability. One of the most significant achievements in that direction is the discovery of Albers-Fuchs-Merry that the unorderability of $\mathcal{G}(M,\xi)$ implies the Weinstein conjecture in $(M,\xi)$ [2]. The goal of this article is to establish an equivalence between orderability and the existence of (hopefully spectral) contact selectors.
@@
    (triangle inequality) $c_{+}(\varphi\psi)\leq c_{+}(\varphi)+c_{+}(\psi)$,
@@ 1.2. Conventions
Whenever no precision regarding regularity is given, maps and manifolds considered are $C^{\infty}$-smooth. Every contact manifold $(M,\xi)$ considered are assumed to be connected and cooriented (i.e. $TM/\xi$ is an oriented line bundle). We write that a contact form $\alpha$ is supporting $\xi$ if $\ker\alpha=\xi$ and $\alpha$ respects the coorientation. A contactomorphism of $(M,\xi)$ will always preserve the coorientation. Given a contact manifold $(M,\xi)$ and a closed Legendrian submanifold $\Lambda_{*}\subset M$, one denotes $\mathcal{L}(\Lambda_{*})$ the isotopy class of $\Lambda_{*}$. One denotes $\mathcal{G}(M)$ the group $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ of compactly supported contactomorphisms of $(M,\xi)$ isotopic to $\mathrm{id}$ through compactly supported contactomorphisms while $\widetilde{\mathcal{G}}(M)$ will denote its universal cover. For more details about the topology considered on these spaces we refer the reader to Section 2.2.
@@
On $O$ being either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, we write $x\preceq y$, or equivalently $y\succeq x$, if there exists a non-negative isotopy in $O$ from $x$ to $y$ (see Section 2 for more details). $O$ is called orderable if and only if $\preceq$ defines a partial order (see Section 2.4 for details). In particular, the orderability of $\mathcal{L}$ (resp. $\mathcal{G}$) implies the orderability of $\widetilde{\mathcal{L}}$ (resp. $\widetilde{\mathcal{G}}$). We refer to Examples 2.14 for a list of known orderable spaces.
@@ 1.3. Order spectral selectors
Let $\Lambda_{*}$ be a closed Legendrian submanifold of $(M,\xi)$ such that $\mathcal{L}(\Lambda_{*})$ (resp. ~$\widetilde{\mathcal{L}}(\Lambda_{*})$) is orderable and let $\alpha$ be a complete contact form supporting $\xi$. The selectors ±$\ell_{\pm}^{\alpha}$ are real-valued and satisfy the following properties for every $\Lambda_{0},\Lambda_{1},\Lambda_{2}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$),
@@
    (compatibility) ±$\ell_{\pm}^{\alpha}(\varphi(\Lambda_{1}),\varphi(\Lambda_{0}))=\ell_{\pm}^{\varphi^{*}\alpha}(\Lambda_{1},\Lambda_{0})$
    (resp. ±$\ell^{\alpha}_{\pm}(\varphi(\Lambda_{1}),\varphi(\Lambda_{2}))=\ell_{\pm}^{\Pi(\varphi)^{*}\alpha}(\Lambda_{1},\Lambda_{0})$), for every $\varphi\in\mathrm{Cont}(M,\xi)$ (resp. $\varphi\in\widetilde{\mathrm{Cont}}(M,\xi)$),
@@
    (non-degeneracy) $\ell_{+}^{\alpha}(\Lambda_{1},\Lambda_{0})=\ell_{-}^{\alpha}(\Lambda_{1},\Lambda_{0})=t$ for some $t\in\mathbb{R}$ implies $\Lambda_{1}=\phi_{t}^{\alpha}\Lambda_{0}$ (resp. $\Pi\Lambda_{1}=\phi_{t}^{\alpha}\Pi\Lambda_{0}\in\mathcal{L}$).
@@
    (compatibility) $c_{\pm}^{\alpha}(\varphi\psi\varphi^{-1})=c_{\pm}^{\varphi^{*}\alpha}(\psi)$ (resp. $c_{\pm}^{\alpha}(\varphi\psi\varphi^{-1})=c_{\pm}^{\Pi(\varphi)^{*}\alpha}(\psi)$), which extends to every $\varphi\in\mathrm{Cont}(M,\xi)$ (resp. $\varphi\in\widetilde{\mathrm{Cont}}(M,\xi)$),
@@
    (non-degeneracy) $c_{-}^{\alpha}(\psi)=c_{+}^{\alpha}(\psi)=t$ for some $t\in\mathbb{R}$ implies $\psi=\phi^{\alpha}_{t}$ (resp. $\Pi\psi=\phi^{\alpha}_{t}\in\mathcal{G}$).
@@
We emphasize the fact that we do not know that the $c_{\pm}^{\alpha}(\psi)$’s do actually select spectral values of $\psi$. Nevertheless, in the few examples we are able to compute explicitly $c^{\alpha}_{\pm}$ (for example see Remark 4.13 and [7, Corollary 5.1]), these maps do take spectral values and therefore strongly suggest that they should be spectral in general.
The following direct corollary states that orderability is equivalent to the existence of selectors in a rather weak sense (a similar statement for spaces $\mathcal{L}$ and $\widetilde{\mathcal{L}}$ also holds). In contrast with their symplectic counterparts, these selectors are not actually invariant: identities $\ell(\varphi\Lambda_{1},\varphi\Lambda_{0})=\ell(\Lambda_{1},\Lambda_{0})$ and $c(\varphi\psi\varphi^{-1})=c(\psi)$ are not verified for every contactomorphism $\varphi$ (resp. lift in the universal cover), $\Lambda_{0},\Lambda_{1}$ and $\psi$ being fixed. In that sense, the $\alpha$-selectors are only invariant under $\alpha$-strict contactomorphisms: ±$\ell^{\alpha}_{\pm}(\varphi\Lambda_{1},\varphi\Lambda_{0})=\ell^{\alpha}_{\pm}(\Lambda_{1},\Lambda_{0})$ and $c^{\alpha}_{\pm}(\varphi\psi\varphi^{-1})=c^{\alpha}_{\pm}(\psi)$ when $\varphi^{*}\alpha=\alpha$. Of course, this is compatible with the fact that the $\alpha$-spectrum is only invariant under strict contactomorphisms: as in general $\varphi^{-1}\phi^{\alpha}_{t}\varphi=\phi^{\varphi^{*}\alpha}_{t}$ for all $t\in\mathbb{R}$, one has
@@
Nonetheless, the sign of these selectors is invariant (see Lemmata 3.2 and 3.8).
@@ 1.4. Applications to the existence of Reeb chords and translated points
[…]
@@
In its original setting, the Arnold chord conjecture asked about the existence of a non-trivial Reeb chord the endpoints of which lie on a same Legendrian circle of the standard contact 3-sphere, given any Legendrian circle [10, §8]. In this setting Mohnke proved a generalization to closed contact manifolds arising as the boundary of subcritical Stein manifolds (Legendrian circles are replaced by closed Legendrian submanifolds) [58]. Hutching and Taubes extended it to any closed contact 3-manifolds [46]. Very recent works still discuss generalized forms of the original conjecture: Chantraine proved a version for a specific kind of fillable contact manifolds with Lagrangian slices taking the place of Legendrian submanifolds [18], the case of Legendrian submanifolds lying in $P\times\mathbb{R}$, for a Liouville manifold $P$, has also received a lot of attention (see [50] and references therein).
@@
Since the discovery by Hofer of a bi-invariant metric on the group of Hamiltonian diffeomorphisms [44], multiple attempts to obtain a contact counterpart where made. On the one hand, the obvious generalization of the Hofer metric, often referred to as the Shelukhin-Hofer metric, is not bi-invariant. On the other hand, one can define interesting bi-invariant metrics on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ but they are discrete, as was discovered by Colin-Sandon [27] and Fraser-Polterovich-Rosen [36]. Our spectral selectors allow us to define a contact analogue of the spectral metric discovered by Viterbo in [75]. We use this new metric to study both Hofer type contact metrics and the discrete bi-invariant metrics of Colin-Sandon and Fraser-Polterovich-Rosen. In particular, this new metric allows us to study questions of non-degeneracy, unboundedness, geodesics and equivalence of the different mentionned metrics. We directly refer to Section 4 for details.
@@ 2.1. Legendrian and contact isotopies and their Hamiltonian maps
Throughout this section $I\subset\mathbb{R}$ will always be an interval containing 0. Let $(M,\xi)$ be a cooriented contact manifold. A contact form $\alpha$ supporting $\xi$ (i.e. $\ker\alpha=\xi$) is called complete if its Reeb vector field $R^{\alpha}$ induces a complete Reeb flow. We recall that the Reeb vector field $R^{\alpha}$ is uniquely defined by the equations $\alpha(R^{\alpha})\equiv 1$ and $\iota_{R^{\alpha}}\mathrm{d}\alpha=0$. Let us fix a complete contact form $\alpha$ supporting $\xi$.
Let us recall that a contact isotopy $(\varphi_{t})_{t\in I}$ is a family of contactomorphisms such that the map $I\times M\to M,(t,x)\mapsto\varphi_{t}(x)$ is smooth. For any contact isotopy $(\varphi_{t})$ such that $\varphi_{0}=\mathrm{id}$, there exists a unique smooth map $H:I\times M\to\mathbb{R}$ related to the time-dependent contact vector field $(X_{t})_{t\in I}$ generating the isotopy, i.e. $\frac{d}{dt}\varphi_{t}=X_{t}\circ\varphi_{t}$, by
@@
where $R^{\alpha}$ denotes the Reeb vector field of $\alpha$ and $H_{t}:M\to\mathbb{R}$ is induced by the restriction of $H$ to $\{t\}\times M$ (see e.g. [37, §2.3]).
Definition 2.1.
Let $(\varphi_{t})_{t\in I}$ be a contact isotopy such that $\varphi_{0}=\mathrm{id}$. We call the map $H:I\times M\to\mathbb{R}$ satisfying (2) the $\alpha$-contact Hamiltonian map of $(\varphi_{t})$.
Conversely, to any smooth map $H:I\times M\to\mathbb{R}$ is associated a unique time-dependent contact vector field $(X_{t})$ by the system (2). When the contact vector field is complete (e.g. when the differential of $H$ is compactly supported), we say that $H$ generates the contact isotopy obtained by integrating $(X_{t})$.
@@
As $j_{t}^{*}\alpha=0$ for all $t\in I$, one can deduce that $(H_{t})$ only depends on the isotopy $(\Lambda_{t})$ and not the specific choice of parametrization $j$.
Definition 2.2.
Let $(\Lambda_{t})_{t\in I}$ be a Legendrian isotopy. The family $(H_{t}:\Lambda_{t}\to\mathbb{R})_{t\in I}$ satisfying (3) is called the $\alpha$-contact Hamiltonian map of $(\Lambda_{t})$.
The contact Hamiltonian map of a Legendrian isotopy $(\Lambda_{t})_{t\in I}$ can be seen as a smooth map $H:N\to\mathbb{R}$ defined on the submanifold $N:=\bigcup_{t}\{t\}\times\Lambda_{t}$ of $I\times M$. If $\Lambda_{t}=\varphi_{t}(\Lambda_{0})$ for some contact isotopy $(\varphi_{t})$ starting at the identity whose $\alpha$-Hamiltonian map is $K:I\times M\to\mathbb{R}$, the Hamiltonian of $(\Lambda_{t})$ is the restriction of $K$ to $N$.
@@
Let $I\subset\mathbb{R}$ be an interval containing 0 and let $(\Lambda_{t})_{t\in I}$ be a Legendrian isotopy on $(M,\ker\alpha)$, the $\alpha$-contact Hamiltonian map of which is $H:\bigcup_{t}\{t\}\times\Lambda_{t}\to\mathbb{R}$. Let $K:I\times M\to\mathbb{R}$ be an $\alpha$-contact Hamiltonian smoothly extending $H$ and generating a contact isotopy $(\varphi_{t})$, then $\varphi_{t}(\Lambda_{0})=\Lambda_{t}$ for all $t\in I$.
Before proving Lemma 2.3, let us recall the link between Legendrian isotopies and the Hamilton-Jacobi equation. Let $(\Lambda_{t})$ be a Legendrian isotopy. According to the Legendrian Weinstein neighborhood theorem, there exists an open subset $W\subset M$ containing $\Lambda_{0}$ that is contactomorphic to a neighborhood of the zero-section of the 1-jet space $J^{1}\Lambda_{0}:=T^{*}\Lambda_{0}\times\mathbb{R}$, through this identification $\Lambda_{0}$ is sent to the zero-section. Moreover, our chosen contact form $\alpha$ is sent to the standard contact form $\alpha_{0}:=\mathrm{d}z-\lambda$, $z$ being the $\mathbb{R}$-coordinate and $\lambda$ being the pull-back of the tautological form of $T^{*}\Lambda_{0}$. With this local identification, for $t\in I$ close to 0, $\Lambda_{t}$ is a Legendrian graph above the zero section $\Lambda_{0}$, so it is the 111-jet of a smooth map $f_{t}:\Lambda_{0}\to\mathbb{R}$,
@@
for all $q\in\Lambda_{0}$ and $t\in I$ close to 0 (we have used that $\alpha$ is identified to $\alpha_{0}$ in $W$).
@@
By connectivity of the interval $I$, it is enough to prove this statement for $t\in I$ close enough to 0. By considering a Weinstein neighborhood of $\Lambda_{0}$ as above, one can then identify $\Lambda_{t}$ and $\varphi_{t}(\Lambda_{0})$ to the 111-jet of the respective maps $f_{t}:\Lambda_{0}\to\mathbb{R}$ and $g_{t}:\Lambda_{0}\to\mathbb{R}$, for all $t\in I$, shrinking $I$ if necessary. Let us remark that the contact Hamiltonian map of $(\varphi_{t}(\Lambda_{0}))$ is the restriction of $K$ to $\bigcup_{t}\{t\}\times\varphi_{t}(\Lambda_{0})$. Therefore, by the above discussion, both families of maps $(f_{t})$ and $(g_{t})$ satisfy the same Hamilton-Jacobi equation
@@
with the initial condition $u_{0}\equiv 0$. By unicity of the smooth solutions of the Hamilton-Jacobi equation in a small interval of time (see e.g. [53, Exercice 3.5.17]), $f_{t}=g_{t}$ for all $t\in I$ and $\Lambda_{t}=\varphi_{t}(\Lambda_{0})$. ∎
@@
Let $(M,\ker\alpha)$ be a contact manifold endowed with a complete contact form $\alpha$. Let $I\subset\mathbb{R}$ be an interval containing 0 and let $(\Lambda_{t})_{t\in I}$ be a (closed) Legendrian isotopy on $(M,\ker\alpha)$, the $\alpha$-contact Hamiltonian map of which is $H$. Then there exist contact Hamiltonians $K,G:I\times M\to\mathbb{R}$ such that $\inf K=\inf H$ and $\sup G=\sup H$, the respective contact flows of which $(\varphi_{t})$ and $(\psi_{t})$ satisfy $\varphi_{t}(\Lambda_{0})=\psi_{t}(\Lambda_{0})=\Lambda_{t}$ for all $t\in I$.
@@
2.2. The $C^{1}$-topology and the universal covers